OpenStudy (anonymous):

Each year Daisy buys a train ticket to visit his aunt for a week. Last summer he paid $135. This year he paid $162. What was the percent increase in the cost of his train ticket?

OpenStudy (y2o2):

increase in price = 162 - 135 = $27 percentage = (27/135)*100 = 20%

OpenStudy (anonymous):

Thanks Y2o2